Female, 77 years old , height 5'2, weight 165lbs

The formula available at the facility is Jevity 1.2.

Product Information:      

Jevity 1.2 contains 237 mL/can                     

Contains 1.2 kcal/mL                       

Protein = 13.2 gm/237 mL       
Fiber = 4.3 gm/237 mL   

FYI---This Jevity product contains 191 mL free water / 237 mL.

Revisit Mrs. Noland’s kcal, protein and fluid needs. What are they?  

Let’s walk through the feeding calculations required. After reviewing the product information and Mrs. Noland’s nutritional needs that you have already sited, answer the following questions:

*determine how many kcals each can of formula would contain. Answer:     

*determine in cans how much formula is required to meet the patient’s energy needs.   Do this by dividing kcal needs by the number of kcal per can. Answer:

*do the same for protein. How many cans of formula are needed to meet the patient’s protein needs?  
Answer #1

Each can contains 237 ml and 1 ml contains 1.2 kcal.hence one can contains 237×1.2= 284.4 kcal..

This patient needs about 1800k cal.. number of cans

   1800/ 284.4= 6.32cans ie 6 to 7 cans.

Proteins needed is 0.8 per gram per kg per day.

0.8× 77=61.6 grams per day.

Total cans needed is 61.6/13.2= 4.66 ie 4 to 5 cans .